OpenStudy (anonymous):

A laboratory has 20 grams of a radioactive substance that decays in such a way that t days later the number of grams left will be y = 20 (1/10)t/15. How many days does it take for half of the substance to decay? A. 4 B. 4.5 C. 9 D. 9.5

OpenStudy (whpalmer4):

\[y = 20(\frac{1}{10})^{t/15}\] To find the value of \(t\) such that \(y(t) = 10\), just plug in the numbers and solve: \[10 = 20(0.1)^{t/15}\]Divide both sides by 20\[0.5=(0.1)^{t/15}\]Remember the property of logarithms that \(\log a^b = b \log a\) and take the log of both sides: \[\log 0.5 = \frac{t}{15} \log 0.1\]I think you should be able to solve it from there.

OpenStudy (whpalmer4):

I find it especially worthwhile to plug the answer back into the original formula to check the answer when doing problems like this!